tickets for a concert were sold to adults for $3 and to students for $2. if the total receipts were $824 and twice as many adult tickets as students were sold, then how many of each were sold?
Found 2 solutions by oscargut, Mathtut:
Answer by oscargut(2103)

number of adult tickets =x
number ofstudents tickets =y
3x+2y=824
2y=x
then
4x=824 then x=206 and y=103

Answer by Mathtut(3670)

a= # of adults s= # of students so 3a+2s=824 and a=2s so substituting a in the second equation ,into the first equation we get 3(2s)+2s=824 so 8s=824
and s=103 so since a=2 times s(103) then a =206
answer: 206 adult tickets and 103 student tickets were sold.